OpenStudy (anonymous):

Multiply (3A+4B)(3A-4B)? ?

OpenStudy (anonymous):

I don't get this at all....

OpenStudy (anonymous):

9a^2-16b^2

OpenStudy (anonymous):

See I got the 9A^2 but I couldn't get the second part.

OpenStudy (anonymous):

you use the FOIL rules. which would be 3a x 3a, then you do 3a times -4b, and 3a times bb. the two 3a x 4b cancle eachother because one is - and one is +. the final part is 4b x -4b which equals 16b^2

OpenStudy (anonymous):

that make more sense?

OpenStudy (anonymous):

\[3A \times 3A + 3A \times -4B +4B \times 3A + 4B \times -4B\] \[=9A ^{2} -12AB +12AB -B ^{2}\] \[=9A ^{2}-4B ^{2}\]

OpenStudy (anonymous):

Yes I think so. I don't really remember learning FOIL. Thank you

OpenStudy (anonymous):

use this method (If you can imagine it in table form with 3A+4B in the column headings, and in the row headings 3A and -4B) sorry can't show the lines. 3A 4B 3A 9A^2 +12AB -4B -12AB -16B^2 now add the boxes 9A^2 +12AB -12AB -16B^2 =9A^2 -16B^2

OpenStudy (anonymous):

it stands for First, Outside, Inside, Last. the order that i multiply the four parts of that sort of problem

OpenStudy (anonymous):

After a while, you get to learn (for this type of question), the rule of the difference of 2 squares. (x+y)(x-y) = x^2 -y^2 here x = 3A y= 4B so you can go immediately to (3A+4B)(3A-4B) = (3A)^2 - (4B)^2 = 9A^2 - 16B^2 Same as before

OpenStudy (anonymous):

I learnt the smiley face method instead of foil both are just as valid Just make sure you end up with four parts before cancelling